OpenStudy (saranya):

In triangle PQR, ∠PQR = 30° and ∠PRQ = 45°. If S is a point on QR such that QS : QR = 2 : 5, then find the ratio of the area of triangle PQS to that of triangle SPR. A. B. C. 3 : 2 D. 2 : 3

OpenStudy (shubhamsrg):

 let M be a point on QR such that PM is perpendicular to QR. area of PQS = 1/2 PM.QS and area of PSR = 1/2 PM.SR thus area(PQS)/area(PSR) = QS/SR or area(PSR)/area(PQS) = SR/QS = (QR - QS)/QS = QR/QS - 1 = 5/2 -1 = 3/2 hence req ans should be 2:3